OpenStudy (anonymous):

integrate x+2/x^2+2x+3

OpenStudy (dumbcow):

complete the square: \[\int\limits \frac{x+2}{(x+1)^2 +2} dx\] u = x+1 du = dx \[\int\limits \frac{u+1}{u^2 +2} du\] split into 2 integrals \[\int\limits \frac{u}{u^2 +2} + \int\limits \frac{1}{u^2 +2}\] first integral: w = u^2 +2 dw = 2u du \[\frac{1}{2} \int\limits \frac{dw}{w} = \frac{1}{2} \ln w\] 2nd integral: u = sqrt2 tan w du = sqrt2 sec^2 w \[\int\limits \frac{1}{2 \sec^2 w} \sqrt{2} \sec^2 w = \frac{\sqrt{2}}{2} w\] plugging back in for x \[\rightarrow \ln (x^2 +2x+3) + \frac{\sqrt{2}}{2} \tan^{-1} (\frac{x+1}{\sqrt{2}}) + C\]